# MULTI-HOP DEEP JOINT SOURCE-CHANNEL CODING WITH DEEP HASH DISTILLATION FOR SEMANTICALLY ALIGNED IMAGE RETRIEVAL

### v1

@@ Abstract
We consider image transmission via deep joint source-channel coding (DeepJSCC) over multi-hop additive white Gaussian noise (AWGN) channels by training a DeepJSCC encoder-decoder pair with a pre-trained deep hash distillation (DHD) module to semantically cluster images, facilitating security-oriented applications through enhanced semantic consistency and improving the perceptual reconstruction quality. We train the DeepJSCC module to both reduce mean square error (MSE) and minimize cosine distance between DHD hashes of source and reconstructed images. Significantly improved perceptual quality as a result of semantic alignment is illustrated for different multi-hop settings, for which classical DeepJSCC may suffer from noise accumulation, measured by the learned perceptual image patch similarity (LPIPS) metric.
@@ 2 Problem Formulation
Consider a source image ${\mathbf{S}}\!\in\!\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of color channels, height, and width, respectively. We aim to wirelessly transmit $\mathbf{S}$} to its destination via $r$ relay nodes $\{\text{R}_{1},\dots,\text{R}_{r}\}$, where adjacent nodes are connected by complex additive white Gaussian noise (AWGN) channels with additive noise terms ${\mathbf{n}}_{i}$, where $i=1,\!\dots,r\!+\!1$. The AWGN components ${\mathbf{n}}_{i}$ are considered to be mutually independent and identically distributed, i.e., we have ${\mathbf{n}}_{1}\sim\!{\mathbf{n}}_{2}\!\sim\!\dots\!\sim{\mathbf{n}}_{r+1}\!\sim\!\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{k})$ for r+1r\!+\!1 hops, where $k$ denotes the number of complex channel symbols. This simplified assumption allows us to gain fundamental insights from the experimental results.
@@
We measure the reconstruction quality with the average peak SNR (PSNR), defined as
$\text{PSNR}=-10\cdot\log_{10}\left(\frac{\|{\mathbf{S}}-\widehat{{\mathbf{S}}}\|^{2}_{2}}{CHW}\right)\text{ dB}.$ (1)
We measure the perceptual quality of reconstructed images $\widehat{{\mathbf{S}}}$ using LPIPS [12], which has been shown to better align with human perception. We next introduce our baseline protocols.
@@ 2.2 Quantize-and-Forward (QF) Protocol
Quantization operations are less complex than decoding operations, which in principle means simpler circuitry and lower total energy consumption [13, Chapter 14.5]. These properties make a QF relaying protocol well-suited for, e.g., relays in remote locations on the edges of core networks or low-latency satellite communications, as they will mainly quantize the received signal and relay it forward through noiseless pipelines obtained by using error correcting codes for each hop, akin to the setup in [7].
Consider an encoder-decoder pair $(f_{Q},d_{Q})$ that is adapted to transmit through an AWGN channel with a fixed SNR. We want to quantize the channel output $\mathbf{y}$} observed at a relay $\text{R}_{1}$ to a bit sequence $\mathbf{b}$} and forward it through a noiseless pipeline (e.g., perfect channel coding) to the destination decoder ($\text{R}_{2}\rightarrow\cdots\rightarrow\text{R}_{r}$). The sequence $\mathbf{b}$} is then dequantized (mapped) to $\widehat{{\mathbf{y}}}$, which is finally used by the decoder $d_{Q}$ to reconstruct the quantized image $\widehat{{\mathbf{S}}}$. We consider the “naive quantization” approach in [7] as a baseline, where we partition ${\mathbf{y}}\in\mathbb{R}^{2k}$ into $2k/N_{Q}$ blocks. Denote $N_{Q}\geq 1$ as the number of real-valued elements of $\mathbf{y}$} per block, and quantize each block with $N_{Q}b$ bits. We compute the centers of the $2^{N_{Q}b}$ codewords with the K-means algorithm [14], and assume that the codebook is available at the relay $\text{R}_{1}$ and the destination. The rate of the vector quantizer is expressed as bits per pixel (bpp), and we compute this rate as $I=\frac{2kN_{Q}b}{N_{Q}HW}$.
@@ 3.3 Experimental Setup
The dataset considered in this work is a subset of the NUS-WIDE dataset [16], consisting of 9,450:1,050:2,1009,\!450\!:\!1,\!050\!:\!2,\!100 training, validation, and test images with $256\times 256$ resolution (${\mathbf{S}}\!\in\!\mathbb{R}^{3\times 256\times 256}$, where exponent corresponds to $C\times H\times W$) and corresponding $N_{\text{cls}}\!=\!21$ dimensional multi-hot encoded class labels $\mathbf{c}$}. We use a pre-trained ResNet50 [17] as the encoder $E_{\theta}$ with the feature dimension $N_{E}\!=\!2048$. We set the hash length as $N_{H}\!=\!64$ bits, and the remaining parameter values are assigned as default in [8].
@@
For DF, we use mini-batch sizes $[20,10,8,6]$ for training $r=[0,1,2,3]$ relays, respectively (larger batch sizes cause memory overflows on NVIDIA Tesla V100 32GB GPUs). We train and test them at SNRs [5,10,15[-5,-10,-15 dB, i.e., we train 12 DF setting models. We present our results in Section 4.1.
@@ 4.1 DF Relaying
Moreover, when any architecture is trained using LPIPS as a loss term, it achieves a lower LPIPS score than our scheme. This is expected, since the training and evaluation metrics coincide. Our architecture optimizes semantic hash alignment, which sacrifices some LPIPS in favor of semantic alignment across multiple hops.
@@ 4.2 QF Relaying
The memory footprint of the full DeepJSCC module (both encoder and decoder) and DHD is $89.2$ MB and $94.4$ MB, respectively. We note that both models’ computational complexities are the same during inference, as only the weights are affected by the different training methodologies while the architectures are the same. We will study the trade-off between architecture complexity and performance, as we estimated empirical time complexity as $0.094$ and $0.046$ seconds per image during training and testing, respectively.
@@ 5 Conclusion
We demonstrated that our multi-hop DeepJSCC-DHD scheme significantly improves semantic alignment between the perceptual quality of the reconstructed images, measured by LPIPS, for both DF and QF relay settings by leveraging semantic clustering via a trained DHD module. Moreover, our experimental results showed that semantic alignment in our scheme remains robust also to quantization effects. Therefore, our multi-hop DeepJSCC-DHD scheme adds a semantic alignment capability to DeepJSCC, complementing perception-oriented training and enabling secure authentication-oriented DeepJSCC applications.
@@ 6 Acknowledgement
This work has been supported by the ZENITH Research and Leadership Career Development Fund under Grant ID23.01, the Swedish Foundation for Strategic Research (SSF) under Grant ID24-0087, and the SNS JU project 6G-GOALS under the EU’s Horizon program (Grant Agreement No. 101139232). The computations and data handling were enabled by resources provided by the National Supercomputer Centre (NSC), funded by Linköping University.

### v2

# MULTI-HOP DEEP JOINT SOURCE-CHANNEL CODING WITH DEEP HASH DISTILLATION FOR SEMANTICALLY ALIGNED IMAGE RECOVERY

@@ Abstract
We consider image transmission via deep joint source-channel coding (DeepJSCC) over multi-hop additive white Gaussian noise (AWGN) channels by training a DeepJSCC encoder-decoder pair with a pre-trained deep hash distillation (DHD) module to semantically cluster images, facilitating security-oriented applications through enhanced semantic consistency and improving the perceptual reconstruction quality. We train the DeepJSCC module to both reduce mean squared error (MSE) and minimize cosine distance between DHD hashes of source and reconstructed images. Significantly improved perceptual quality as a result of semantic alignment is illustrated for different multi-hop settings, for which classical DeepJSCC may suffer from noise accumulation, measured by the learned perceptual image patch similarity (LPIPS) metric.
@@ 2 Problem Formulation
Consider a source image ${\mathbf{S}}\!\in\!\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of color channels, height, and width, respectively. We aim to wirelessly transmit $\mathbf{S}$} to its destination via $r$ relay nodes $\{\text{R}_{1},\dots,\text{R}_{r}\}$, where adjacent nodes are connected by complex additive white Gaussian noise (AWGN) channels with additive noise terms ${\mathbf{n}}_{i}$, where $i=1,\!\dots,r\!+\!1$. The AWGN components ${\mathbf{n}}_{i}$ are considered to be mutually independent and identically distributed, i.e., we have ${\mathbf{n}}_{1}\sim\!{\mathbf{n}}_{2}\!\sim\!\dots\!\sim{\mathbf{n}}_{r+1}\!\sim\!\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{k})$ for r+1r\!+\!1 hops, where $k$ denotes the number of complex channel symbols. This simplified model allows us to gain fundamental insights from the experimental results.
@@
We measure the reconstruction quality of reconstructed images $\widehat{{\mathbf{S}}}$ with the average peak SNR (PSNR), defined as
$\text{PSNR}=-10\cdot\log_{10}\left(\frac{\|{\mathbf{S}}-\widehat{{\mathbf{S}}}\|^{2}_{2}}{CHW}\right)\text{ dB}$ (1)
and the perceptual quality with LPIPS [19], which has been shown to better align with human perception. We next introduce our baseline protocols.
@@ 2.2 Quantize-and-Forward (QF) Protocol
Quantization operations are less complex than decoding operations, which in principle means simpler circuitry and lower total energy consumption [12, Chapter 14.5]. These properties make a QF relaying protocol well-suited for, e.g., relays in remote locations on the edges of core networks or low-latency satellite communications, as they will mainly quantize the received signal and relay it forward through noiseless pipelines obtained by using error correcting codes for each hop, akin to the setup in [3].
Consider an encoder-decoder pair $(f_{Q},d_{Q})$ that is adapted to transmit through an AWGN channel with a fixed SNR. We want to quantize the channel output $\mathbf{y}$} observed at relay $\text{R}_{1}$ to a bit sequence $\mathbf{b}$} and forward it through a noiseless pipeline (e.g., perfect channel coding) to the destination decoder ($\text{R}_{2}\rightarrow\cdots\rightarrow\text{R}_{r}$). The sequence $\mathbf{b}$} is then dequantized (mapped) to $\widehat{{\mathbf{y}}}$, which is finally used by the decoder $d_{Q}$ to reconstruct the quantized image $\widehat{{\mathbf{S}}}$. We consider the “naive quantization” approach in [3] as a baseline, where we partition ${\mathbf{y}}\in\mathbb{R}^{2k}$ into $2k/N_{Q}$ blocks. Denote $N_{Q}\geq 1$ as the number of real-valued elements of $\mathbf{y}$} per block, and quantize each block with $N_{Q}b$ bits. We compute the centers of the $2^{N_{Q}b}$ codewords with the K-means algorithm [13], and assume that the codebook is available at the relay $\text{R}_{1}$ and the destination. The rate of the vector quantizer is expressed as bits per pixel (bpp), and we compute this rate as $I=\frac{2kN_{Q}b}{N_{Q}HW}$.
@@ 3.3 Experimental Setup
The dataset considered in this work is a subset of the NUS-WIDE dataset [6], consisting of 9,450:1,050:2,1009,\!450\!:\!1,\!050\!:\!2,\!100 training, validation, and test images with $256\times 256$ resolution (${\mathbf{S}}\!\in\!\mathbb{R}^{3\times 256\times 256}$, where exponent corresponds to $C\times H\times W$) and corresponding $N_{\text{cls}}\!=\!21$ dimensional multi-hot encoded class labels $\mathbf{c}$}. We use a pre-trained ResNet50 [9] as the encoder $E_{\theta}$ with the feature dimension $N_{E}\!=\!2048$. We set the hash length as $N_{H}\!=\!64$ bits, and the remaining parameter values are assigned as default in [10]. The weights of $H_{\theta}$ and $E_{\theta}$ are then trained/fine-tuned as in [10] and frozen.
@@
For DF, we use mini-batch sizes $[20,10,8,6]$ for training $r=[0,1,2,3]$ relays, respectively (larger batch sizes cause memory overflows on NVIDIA Tesla V100 32GB GPUs). We train and test them at SNRs [5,10,15][-5,-10,-15] dB, i.e., we train 12 DF setting models. We present our results in Section 4.1.
@@ 4.1 DF Relaying
Training all architectures in 10-10 dB DF scenario using LPIPS with weight $\alpha\in[0.01,1]$ as an additional loss term, we observe that for all $\alpha$ our proposed architecture improves semantic alignment, measured by (2), compared to the baseline. As $\alpha$ increases, PSNR and LPIPS values of our architecture converge toward those of the baseline, with roughly equivalent PSNR and LPIPS values when training with $\alpha\geq 0.6$. Thus, our architecture significantly improves semantic-alignment performance when trained with LPIPS, with similar perceptual quality to the baseline.
@@ 4.2 QF Relaying
The memory footprints of the full DeepJSCC module (both encoder and decoder) and DHD are $89.2$ MB and $94.4$ MB, respectively. We note that both models’ computational complexities are the same during inference, as only the weights are affected by the different training methodologies while the architectures are the same. We will study the trade-off between architecture complexity and performance, as we estimated empirical time complexity as $0.094$ and $0.046$ seconds per image during training and testing, respectively.
@@ 5 Conclusion
We demonstrated that our multi-hop DeepJSCC-DHD scheme significantly improves semantic alignment between the perceptual quality of the reconstructed images, measured by LPIPS, for both DF and QF relay settings by leveraging semantic clustering via a trained DHD module. Moreover, our experimental results showed that semantic alignment in our scheme remains robust also to quantization effects. Therefore, the proposed multi-hop DeepJSCC-DHD scheme adds a semantic alignment capability to DeepJSCC, complementing perception-oriented training and enabling secure authentication-oriented DeepJSCC applications.
@@ 6 Acknowledgement
This work has been supported by the ZENITH Research and Leadership Career Development Fund under Grant ID23.01, the Swedish Foundation for Strategic Research (SSF) under Grant ID24-0087, the SNS JU project 6G-GOALS under the EU’s Horizon program (Grant Agreement No. 101139232), and the German Federal Ministry of Research, Technology and Space (BMFTR) 6GEM+ Transfer Hub under Grant 16KIS2412. The computations and data handling were enabled by resources provided by the National Supercomputer Centre (NSC), funded by Linköping University.